OpenStudy (anonymous):

Twice the difference between a number and 4 is twenty. Find the number. 8 9 12 14

OpenStudy (anonymous):

I believe it is 14. TWICE the number is twenty, which means the number is ten.

OpenStudy (anonymous):

Add ten to 4... get 14 :)

OpenStudy (anonymous):

\(20=2(x-4)\) \(20=2x-8\) \(28=2x\) \(x=14\) You are right, @roseflutterwings :)

OpenStudy (anonymous):

Yay! Got that, @deeeee ?